# Simultaneously search for multi-target Galactic binary gravitational waves in reduced parameter space with LMPSO-CV

### v1

@@ IV.1 Data preprocessing
Before initiating the search, it is essential to preprocess the LDC1-4 data. Assuming conventional methods have detected signals with SNR $>$ 15, our initial step involves subtracting all injected Galactic binary signals with optimal SNR $>$ 15 from the raw data, allowing us to focus on the search for lower SNR signals. Then the data undergoes band-pass filtering and is divided into 1491 bins for parallel computation, covering a frequency range from 0.09 mHz to 15 mHz. The band-pass filter method aligns with the one illustrated in Figure 3 of Ref. [38], and other techniques, such as undersampling to reduce data length, follow the approach outlined in the same reference. As our method avoids subtracting any signal during the search, concerns about inaccurate subtraction contamination are minimized. Our primary concern shifts to addressing the potential presence of degeneracy noise at the edges of each bin. In preparation for subsequent degeneracy noise removal steps, we slightly expand the frequency search range on both sides of the acceptance zone. This extension encompasses 0.025% times outward from the boundary of the acceptance zone, ensuring coverage of all signals outside the acceptance zone whose degeneracy noise may affect the acceptance zone. This expansion is visually represented in Fig. 11.
@@ IV.2 Local maxima PSO
The Particle Swarm Optimization (PSO) algorithm, renowned for its efficacy in seeking global maxima of continuous nonlinear functions [51, 52], has proven instrumental in the quest for Galactic binaries [38]. Previous methodologies focused on identifying global maxima of the $\mathcal{F}$-statistic within parameter space, employing multiple searches to ensure the robustness of global maxima detection. This process involved iteratively subtracting the detected signal and initiating subsequent search rounds, known as the iterative-subtraction strategy. In our LMPSO-CV approach, we streamline the process by concentrating solely on searching for local maxima, enabling us to achieve convergence with fewer steps. To facilitate this, we have tailored the PSO algorithm to swiftly converge on local maxima, optimizing the efficiency of the search process. We employ a total of 40 particles in our algorithm. The process begins by randomly dispersing these particles throughout the parameter space. Each particle calculates the $\mathcal{F}$-statistic value at its respective position. Based on the calculation results, the positions of the particles undergo iteration. Each particle “remembers” the best value and the position it explored, denoted as the personal best position ($P$). From the $P$ of all particles, the global best position ($G$) is determined to represent the globally best position found by the swarm. To reach the optimal solution, each particle moves toward both its $P$ and the $G$ within the swarm. The position of each particle is updated in each iteration through the following equation:
@@
In the given equation, $t$ denotes the number of iterations. $\mathbf{v}$ represents the velocity vector. $\omega$ serves as the inertia weight crucial for balancing local exploitation and global exploration. $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ denote random vectors uniformly distributed within the range $[0,1]^{D}$ ($D$ is the dimensionality of the searched parameter space). $\mathbf{x}_{i}$ indicates the position of the $i$-th particle. $c_{1}$ and $c_{2}$, referred to as “acceleration coefficients”, are positive constants. In particular, when $c_{1}=c_{2}=2$ and $\omega=\text{max}(\omega_{1}-\frac{\omega_{2}}{\omega_{3}}t,\leavevmode\nobreak\ \omega_{4})$ (where $\omega_{1}=0.9,\leavevmode\nobreak\ \omega_{2}=0.4,\leavevmode\nobreak\ \omega_{3}=1999,\leavevmode\nobreak\ \omega_{4}=0.2$), this parameter configuration strikes a harmonious balance between $P$ and $G$, rendering the PSO algorithm highly effective in the pursuit of the global maxima. In the employed PSO algorithm, we initially configure the parameters as described above. Upon reaching an optimal $\mathcal{F}$-statistic surpassing a predefined threshold ($\mathcal{F}$-statistic $>$ 53, corresponding to optimal SNR $>$ 7), we dynamically adjust the parameter configuration by setting $c_{1}=\omega=0$ in the iterative equation. So the iterative equation becomes:
@@
This adjustment accelerates the convergence of the PSO algorithm towards a local maxima. To ascertain convergence, we implement an automatic convergence criterion: if the optimal $\mathcal{F}$-statistic value remains constant for 15 consecutive iterations, convergence is determined, and facilitating the transition to search the next local maxima. The flow chart of LMPSO in Fig. 14 illustrates the workflow of the LMPSO algorithm. Following the completion of each local maxima search, we create a void in the parameter space corresponding to the position of the $G$. In the subsequent search, particles are initially assessed to determine if they fall within the voids. If a particle lands in the voids, its value is assigned negative infinity, indicating that the particle is invalid. For the algorithm, this signifies a position that requires no further exploration. During the next iteration, the particle will navigate to other positions in the parameter space. As the maximum $\mathcal{F}$-statistic value surpasses a predefined threshold ($\mathcal{F}$-statistic $>$ 53), the search transitions from global to local maxima. Ultimately, the convergence criterion adjudicates whether the search is concluded.
@@ IV.3 Create voids in the parameter space
After identifying a local maxima using LMPSO algorithm, we create a void in the parameter space to prevent redundant searches. The void is modeled as a spheroid described by the ellipsoid equation ($\frac{x^{2}}{a^{2}}+\frac{y^{2}}{b^{2}}+\frac{z^{2}}{c^{2}}=1$), where $a$, $b$, and $c$ represent the radii of the spheroid in $f$, $\beta$, and $\lambda$, respectively. We set $a=df$ to be equal to the minimum frequency resolution. The values for $b$ and $c$ depend on the $\mathcal{F}$-statistic degeneracy calculated in Fig. 12. In this figure, while keeping the other parameters constant, we vary one parameter at a time to observe the trend of the $\mathcal{F}$-statistic in $\beta$ or $\lambda$. Specifically, Fig. 12(a) explores the variation with respect to $\beta$, and Fig. 12(b) with respect to $\lambda$. The resolution of $\beta$ and $\lambda$ is $\pi/40$ radian. We determine the ellipsoidal radius by selecting half the distance between the two local minima closest to the searched local maxima. The positions of these two local minima are denoted by the left boundary and right boundary lines in Fig. 12(a) and Fig. 12(b), with the “true-F” marking the real binary position.
Figure 12: We conduct an analysis of the normalized $\mathcal{F}$-statistic degeneracy to determine the void radii within the Ecliptic Latitude $\beta$ and Ecliptic Longitude $\lambda$. Throughout this computation, all other parameters are kept constant while one parameter is systematically varied at a time. Specifically, in Fig. 12(a), we investigate the effects of changing $\beta$, and in Fig. 12(b), we explore the impact of varying $\lambda$. The spatial resolution employed is $\pi/40$ radians. The determination of the ellipsoidal radius associated with each parameter involves selecting half the distance between the two local minima, identified by the left boundary and right boundary, which are closest to the sought local maxima. The position of the sought local maxima is denoted by the term “true-F”, symbolizing the authentic binary position in our analysis.
@@ V.1 Remove individual signal degeneracy noise
The degeneracy noise produced by an individual signal is the most numerous and easiest to remove, characterized by relatively weaker $\mathcal{F}$-statistic values compared to the binaries from which they originate, so we commence by systematically eliminating individual signal degeneracy noise. The starting point for this removal process is the descending list of searched local maxima, ordered by their $\mathcal{F}$-statistic values. Beginning with the local maxima possessing the highest $\mathcal{F}$-statistic value, indicative of potential authenticity as a real binary, we generate its LISA response GW signal as the initial waveform. Having obtained the initial waveform, we proceed to scrutinize the second-largest $\mathcal{F}$-statistic value to ascertain whether it corresponds to the degeneracy noise of the first. This involves computing the $\mathcal{F}$-statistic value using the parameters of the second local maxima on the first waveform. Subsequently, we deduct the computed $\mathcal{F}$-statistic value of the second local maxima on the first waveform from its own searched $\mathcal{F}$-statistic value. If the difference exceeds the threshold ($\Delta\mathcal{F}$-statistic $>$ 53), it is considered a new signal to retain. We then generate its waveform and incorporate it into the first waveform. Conversely, if the difference falls below the threshold, it is identified as noise and promptly removed. The formula for computing the $\Delta\mathcal{F}$-statistic is given by:
@@
where $S$ represents the raw data, specifically the LDC1-4 residual data in this article. The term $\mathcal{F}\left(\theta_{i},S\right)$ denotes the $\mathcal{F}$-statistic value for $i$ local maxima with parameters $\theta_{i}$ concerning the data $S$ (the searched $\mathcal{F}$-statistic value). Similarly, $\mathcal{F}\left(\theta_{i},\sum\limits_{m=1}^{i-1}h\left(\theta_{m}\right)\right)$ represents the $\mathcal{F}$-statistic value for $i$ local maxima concerning the sum of the first $i-1$ retained waveforms (the computed $\mathcal{F}$-statistic value). This iterative procedure continues, processing all local maxima in descending order of $\mathcal{F}$-statistic. Ultimately, this step significantly reduces the number of initially searched local maxima from the total count of 646,120 (646,120 with $\mathcal{F}$-statistic $>$ 53, and the overall total is 849,576) to 25,615, which are the ones remaining within the acceptance zone after the removal of individual signal degeneracy noise. The flow chart of degeneracy noise removal in Fig. 14 visually outlines the process.
@@ V.2.1 The second search based on astrophysical models
Considering that the majority of Galactic binaries involve white dwarfs with masses ranging from $0.1M_{\odot}$ [53] to $1.4M_{\odot}$, we determine the range of the chirp mass ${\cal M}=(m_{1}m_{2})^{3/5}/M^{1/5}$. Here, $m_{1}$ and $m_{2}$ represent the masses of the two white dwarfs, and $M=m_{1}+m_{2}$ denotes their total mass. The lower and upper limits of the chirp mass ${\cal M}$ are ${\cal M}_{\text{min}}=0.06M_{\odot}$ and ${\cal M}_{\text{max}}=1.22M_{\odot}$, respectively. In contrast to the first search, the $\dot{f}$ range for the second search is constrained to a narrower interval, with its magnitude being proportional to the frequency.
@@ V.2.3 Restrict the Galactic Latitude
The distribution of degeneracy noise spans the entire celestial sphere. However, leveraging the well-established concentration of Galactic binaries in the Galactic Disk, we can enhance the precision of the search results by introducing constraints on their sky locations. By confining the Galactic Latitude of the outcomes within the range of -0.5 rad to 0.5 rad (as illustrated in Fig. 1), we effectively narrow down the selection to 9,143 remaining local maxima.
V.3 Remove overlapped signals degeneracy noise
In the final phase, our aim is to eliminate the overlapped degeneracy noise originating from the simultaneous presence of two or more binaries (as depicted in Fig. 7(c)). Since the $\mathcal{F}$-statistic value of such noise exceeds that of any individual binary it emanates from, it cannot be effectively addressed in the initial step designed to remove individual signal degeneracy noise. To tackle these instances of overlapped degeneracy noise, we repeat the first step in ascending order of the searched local maxima, as illustrated in the flow chart of degeneracy noise removal depicted in Fig. 14. The waveform is initially generated by the signal with the lowest $\mathcal{F}$-statistic value and progressively incorporates signals with higher $\mathcal{F}$-statistic value. The objective is to discern whether the signal with the higher $\mathcal{F}$-statistic value is degeneracy noise resulting from overlapping with signals with lower $\mathcal{F}$-statistic values. Following this step, 8,995 local maxima persist.
@@ VI Search results
where $\theta$ and $\theta^{\prime}$ represent the parameters, and $I\in\{A,E,T\}$. We selected injection sources with 3 $<$ optimal SNR $<$ 15 over a two-year observation period for calculation, with a total count of 55,714. A source is considered accurate if its $R$geq 0.9, relatively accurate if 0.8 $\leq$ $R$ $<$ 0.9, and contaminated if 0.5 $\leq$ $R$ $<$ 0.8.
VI.1 False source fraction and SNR
@@
We employ the reported sources ($N_{r}$) to represent the outcomes following the $-\mathcal{F}$-statistic-analysis, and confirmed sources ($N_{c}$) to signify those reported sources with an $R$ value exceeding the established threshold. The false source fraction (FAS) is succinctly formulated as follows:
$\text{FAS}=1-\frac{N_{c}}{N_{r}}.$ (24)
@@
This metric quantifies the ratio of sources with $R$ values that fail to meet the specified threshold. The evolutionary trajectory of our methodology through the various stages of $-\mathcal{F}$-statistic-analysis (elaborated in Sec. V) is vividly depicted in the histogram of $R$ presented in Fig. 15. In the initial step, a significant reduction in individual degeneracy noise was achieved, resulting in 25,615 reported sources with a FAS of 76.8% at $R$geq 0.8 (A). The second step concentrated on diminishing instrument noise through cross-validation with a secondary search, yielding 10,623 reported sources and a FAS of 54.8% at $R$geq 0.8 (B2). Subsequent refinement in the third step, which confined Galactic Latitude within ±$\pm 0.5$ rad, led to a reduction to 9,143 reported sources and a FAS of 48.0% at $R$geq 0.8 (B3). Finally, the fourth step addressed overlapped degeneracy noise with $\mathcal{F}$-statistic values surpassing those of any individual signal, resulting in 8,995 reported sources and a FAS of 47.7% at $R$geq 0.8 (C). In Fig. 15, the designation “D” characterizes the capability to identify reliable sources based on the variation in $\mathcal{F}$-statistic values before and after the $-\mathcal{F}$-statistic-analysis. Specifically, we select sources for which the $\mathcal{F}$-statistic value demonstrates a change of less than 10% before and after the $-\mathcal{F}$-statistic-analysis, resulting in a FAS of 26.9% (D). However, it is crucial to note that this method is effective mainly for selecting more isolated sources unaffected by other signals. Consequently, the number of reported sources decreases to 3,463. Therefore, we do not consider this approach as an integral part of the $-\mathcal{F}$-statistic-analysis. In the subsequent sections of the article, we exclusively present the reported sources obtained through the $-\mathcal{F}$-statistic-analysis, rather than the “D” result with a lower false source fraction.
@@
Figure 15: The histogram depicts the distribution of correlation coefficients in the results after each step. A, B2, B3, and C correspond to the four steps outlined in Sec. V. A signifies the outcomes after eliminating individual signal degeneracy noise (Sec. V.1). B2 represents the results after cross-validation with the second search (Sec. V.2.2). B3 indicates the outcomes after restricting the Galactic Latitude (Sec. V.2.3), and C denotes the reported sources after removing overlapped degeneracy noise (Sec. V.3). D designates the sources for which the change in $\mathcal{F}$-statistic value is less than 10% before and after the identification of the real $\mathcal{F}$-statistic analysis.
@@
Owing to signal overlap, the $\mathcal{F}$-statistic value for each source may receive contributions from other sources, resulting in an inflated SNR. To illustrate this effect, we depict the optimal SNR ratio between confirmed sources (with $R$geq 0.8) and their corresponding injection sources in Fig. 16.
SNR $\text{SNR ratio}=\frac{\text{SNR}_{\text{{con}}}}{\text{SNR}_{\text{{inj}}}}.$ (25)
@@
The sources are arranged in ascending frequency order, and to achieve a smoother curve, we take an average for every 50 sources. In Fig. 16, “unremove degeneracy noise” refers to the SNR before the $-\mathcal{F}$-statistic analysis, while “remove degeneracy noise” corresponds to the SNR after the $-\mathcal{F}$-statistic analysis (the steps for removing degeneracy noise are outlined in Sec. V.1 and Sec. V.3). The average SNR ratio is 1.22 and 1.10 for “unremove degeneracy noise” and “remove degeneracy noise”, respectively. After the $-\mathcal{F}$-statistic analysis, the fact that the average SNR ratio approaches 1 indicates a closer match between the SNR of the confirmed sources and their corresponding injection sources, as the $\mathcal{F}$-statistic component from other sources has been eliminated. As the frequency increases, the SNR ratio gradually converges to 1. However, at lower frequencies ($f<2\times 10^{-3}$ Hz), a substantial number of unresolved Galactic binaries contribute energy (degeneracy noise) to the confirmed sources, leading to $\text{SNR}_{\text{{con}}}$ being significantly larger than their corresponding $\text{SNR}_{\text{{inj}}}$.

### v3

@@ IV.1 Data preprocessing
Before initiating the search, it is essential to preprocess the LDC1-4 data. Assuming conventional methods have detected signals with SNR $>$ 15, our initial step involves subtracting all injected Galactic binary signals with SNR $>$ 15 from the raw data, allowing us to focus on the search for lower SNR signals. Then the data undergoes band-pass filtering and is divided into 1491 bins for parallel computation, covering a frequency range from 0.09 mHz to 15 mHz. The band-pass filter method aligns with the one illustrated in Figure 3 of Ref. [38], and other techniques, such as undersampling to reduce data length, follow the approach outlined in the same reference. As our method avoids subtracting any signal during the search, concerns about inaccurate subtraction contamination are minimized. Our primary concern shifts to addressing the potential presence of degeneracy noise at the edges of each bin. In preparation for subsequent degeneracy noise removal steps, we slightly expand the frequency search range on both sides of the acceptance zone. This extension encompasses 0.025% times outward from the boundary of the acceptance zone, ensuring coverage of all signals outside the acceptance zone whose degeneracy noise may affect the acceptance zone. This expansion is visually represented in Fig. 10.
@@ IV.2 Local maxima PSO
The Particle Swarm Optimization (PSO) algorithm, renowned for its efficacy in seeking global maxima of continuous nonlinear functions [53, 54], has proven instrumental in the quest for Galactic binaries [38]. Previous methodologies focused on identifying global maxima of the $\mathcal{F}$-statistic within parameter space, employing multiple searches to ensure the robustness of global maxima detection. This process involved iteratively subtracting the detected signal and initiating subsequent search rounds, known as the iterative-subtraction strategy. In our LMPSO-CV approach, we streamline the process by concentrating solely on searching for local maxima, enabling us to achieve convergence with fewer steps. To facilitate this, we have tailored the PSO algorithm to swiftly converge on local maxima, optimizing the efficiency of the search process.
We employ a total of 40 particles in our algorithm. The process begins by randomly dispersing these particles throughout the parameter space. Each particle calculates the $\mathcal{F}$-statistic value at its respective position. Based on the calculation results, the positions of the particles undergo iteration. Each particle “remembers” the best value and the position it explored, denoted as the personal best position ($P$). From the $P$ of all particles, the global best position ($G$) is determined to represent the globally best position found by the swarm. To reach the optimal solution, each particle moves toward both its $P$ and the $G$ within the swarm. The position of each particle is updated in each iteration through the following equation:
@@
In the given equation, $t$ denotes the number of iterations. $\mathbf{v}$ represents the velocity vector. $\omega$ serves as the inertia weight crucial for balancing local exploitation and global exploration. $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ denote random vectors uniformly distributed within the range $[0,1]^{D}$ ($D$ is the dimensionality of the searched parameter space). $\mathbf{x}_{i}$ indicates the position of the $i$-th particle. $c_{1}$ and $c_{2}$, referred to as “acceleration coefficients”, are positive constants. In particular, when $c_{1}=c_{2}=2$ and $\omega=\text{max}(\omega_{1}-\frac{\omega_{2}}{\omega_{3}}t,\leavevmode\nobreak\ \omega_{4})$ (where $\omega_{1}=0.9,\leavevmode\nobreak\ \omega_{2}=0.4,\leavevmode\nobreak\ \omega_{3}=1999,\leavevmode\nobreak\ \omega_{4}=0.2$), this parameter configuration strikes a harmonious balance between $P$ and $G$, rendering the PSO algorithm highly effective in the pursuit of the global maxima.
In the employed PSO algorithm, we initially configure the parameters as described above. Upon reaching an optimal $\mathcal{F}$-statistic surpassing a predefined threshold ($\mathcal{F}$-statistic $>$ 53, corresponding to SNR $>$ 7), we dynamically adjust the parameter configuration by setting $c_{1}=\omega=0$ in the iterative equation. So the iterative equation becomes:
@@
This adjustment accelerates the convergence of the PSO algorithm towards a local maxima. To ascertain convergence, we implement an automatic convergence criterion: if the optimal $\mathcal{F}$-statistic value remains constant for 15 consecutive iterations, convergence is determined, and facilitating the transition to search the next local maxima.
The flow chart of LMPSO in Fig. 13 illustrates the workflow of the LMPSO algorithm. Following the completion of each local maxima search, we create a void in the parameter space corresponding to the position of the $G$. In the subsequent search, particles are initially assessed to determine if they fall within the voids. If a particle lands in the voids, its value is assigned negative infinity, indicating that the particle is invalid. For the algorithm, this signifies a position that requires no further exploration. During the next iteration, the particle will navigate to other positions in the parameter space. As the maximum $\mathcal{F}$-statistic value surpasses a predefined threshold ($\mathcal{F}$-statistic $>$ 53), the search transitions from global to local maxima. Ultimately, the convergence criterion adjudicates whether the search is concluded.
@@ IV.3 Create voids in the parameter space
After identifying a local maxima using LMPSO algorithm, we create a void in the parameter space to prevent redundant searches. The void is modeled as a spheroid described by the ellipsoid equation ($\frac{x^{2}}{a^{2}}+\frac{y^{2}}{b^{2}}+\frac{z^{2}}{c^{2}}=1$), where $a$, $b$, and $c$ represent the radii of the spheroid in $f$, $\beta$, and $\lambda$, respectively. We set $a=df$ to be equal to the minimum frequency resolution. The values for $b$ and $c$ depend on the parameter degeneracy calculated in Fig. 11. In this figure, while keeping the other parameters constant, we vary one parameter at a time to observe the trend of the $\mathcal{F}$-statistic in $\beta$ or $\lambda$. Specifically, Fig. 11(a) explores the variation with respect to $\beta$, and Fig. 11(b) with respect to $\lambda$. The resolution of $\beta$ and $\lambda$ is $\pi/40$ radian. We determine the ellipsoidal radius by selecting half the distance between the two local minima closest to the searched local maxima. The positions of these two local minima are denoted by the left boundary and right boundary lines in Fig. 11(a) and Fig. 11(b), with the red cross marking the real binary position.
Figure 11: We conduct an analysis of the normalized parameter degeneracy to determine the void radii within the Ecliptic Latitude $\beta$ and Ecliptic Longitude $\lambda$. Throughout this computation, all other parameters are kept constant while one parameter is systematically varied at a time. Specifically, in Fig. 11(a), we investigate the effects of changing $\beta$, and in Fig. 11(b), we explore the impact of varying $\lambda$. The spatial resolution employed is $\pi/40$ radians. The determination of the ellipsoidal radius associated with each parameter involves selecting half the distance between the two local minima, identified by the left boundary and right boundary, which are closest to the sought local maxima. The position of the sought local maxima is denoted by the red cross, symbolizing the authentic binary position in our analysis.
@@ V.1 Remove individual signal degeneracy noise
The degeneracy noise produced by an individual signal is the most numerous and easiest to remove, characterized by relatively weaker $\mathcal{F}$-statistic values compared to the binaries from which they originate, so we commence by systematically eliminating individual signal degeneracy noise. The starting point for this removal process is the descending list of searched local maxima, ordered by their $\mathcal{F}$-statistic values. Beginning with the local maxima possessing the highest $\mathcal{F}$-statistic value, indicative of potential authenticity as a real binary, we generate its LISA response GW signal as the initial waveform. Having obtained the initial waveform, we proceed to scrutinize the second-largest $\mathcal{F}$-statistic value to ascertain whether it corresponds to the degeneracy noise of the first. This involves computing the $\mathcal{F}$-statistic value using the parameters of the second local maxima on the first waveform. Subsequently, we deduct the computed $\mathcal{F}$-statistic value of the second local maxima on the first waveform from its own searched $\mathcal{F}$-statistic value. If the residue exceeds the threshold ($\Delta\mathcal{F}$-statistic $>$ 53), it is considered a new signal to retain. We then generate its waveform and incorporate it into the first waveform. Conversely, if the residue falls below the threshold, it is identified as noise and promptly removed. The formula for computing the $\Delta\mathcal{F}$-statistic is given by:
@@
where $S$ represents the raw data, specifically the LDC1-4 residual data in this article. The term $\mathcal{F}\left(\theta_{i},S\right)$ denotes the $\mathcal{F}$-statistic value for $i$ local maxima with parameters $\theta_{i}$ concerning the data $S$ (the searched $\mathcal{F}$-statistic value). Similarly, $\mathcal{F}\left(\theta_{i},\sum\limits_{m=1}^{i-1}h\left(\theta_{m}\right)\right)$ represents the $\mathcal{F}$-statistic value for $i$ local maxima concerning the sum of the first $i-1$ retained waveforms (the computed $\mathcal{F}$-statistic value).
This iterative procedure continues, processing all local maxima in descending order of $\mathcal{F}$-statistic. Ultimately, this step significantly reduces the number of initially searched local maxima from the total count of 646,120 (646,120 with $\mathcal{F}$-statistic $>$ 53, and the overall total is 849,576) to 25,615, which are the ones remaining within the acceptance zone after the removal of individual signal degeneracy noise. The flow chart of degeneracy noise removal in Fig. 13 visually outlines the process.
@@ V.2.1 The second search based on astrophysical models
Considering that the majority of Galactic binaries involve white dwarfs with masses ranging from $0.1\leavevmode\nobreak\ M_{\odot}$ [55] to $1.4\leavevmode\nobreak\ M_{\odot}$, we determine the range of the chirp mass $\mathcal{M}=(m_{1}m_{2})^{3/5}/M^{1/5}$. Here, $m_{1}$ and $m_{2}$ represent the masses of the two white dwarfs, and $M=m_{1}+m_{2}$ denotes their total mass. The lower and upper limits of the chirp mass $\mathcal{M}$ are $\mathcal{M}_{\text{min}}=0.06\leavevmode\nobreak\ M_{\odot}$ and $\mathcal{M}_{\text{max}}=1.22\leavevmode\nobreak\ M_{\odot}$, respectively. In contrast to the first search, the $\dot{f}$ range for the second search is constrained to a narrower interval, with its magnitude being proportional to the frequency.
@@ V.2.3 Restrict the Galactic Latitude
The distribution of degeneracy noise spans the entire celestial sphere. However, leveraging the well-established concentration of Galactic binaries in the Galactic Disk, we can enhance the precision of the search results by introducing constraints on their sky locations. By confining the Galactic Latitude of the outcomes within the range of --0.5 rad to 0.5 rad (as illustrated in Fig. 1), we effectively narrow down the selection to 9,143 remaining local maxima.
V.3 Remove overlapping signals degeneracy noise
In the final phase, our aim is to eliminate the overlapping degeneracy noise originating from the simultaneous presence of two or more binaries (as depicted in Fig. 7(c)). Since the $\mathcal{F}$-statistic value of such noise exceeds that of any individual binary it emanates from, it cannot be effectively addressed in the initial step designed to remove individual signal degeneracy noise. To tackle these instances of overlapping degeneracy noise, we take out the sources for which the difference of the $\mathcal{F}$-statistic values before and after the first step is less than 30% (Sec. V.1) and repeat the first step in ascending order of the searched local maxima, as illustrated in the flow chart of degeneracy noise removal depicted in Fig. 13. The waveform is initially generated by the signal with the lowest $\mathcal{F}$-statistic value and progressively incorporates signals with higher $\mathcal{F}$-statistic value. The objective is to discern whether the signal with the higher $\mathcal{F}$-statistic value is degeneracy noise resulting from overlapping with signals with lower $\mathcal{F}$-statistic values.
Following this step, 6,508 local maxima persist.
@@ VI Search results
where $\theta$ and $\theta^{\prime}$ represent the parameters, and $I\in\{A,E,T\}$. We selected injection sources with 3 $<$ SNR $<$ 15 over a two-year observation period for calculation, with a total count of 55,714. A source is considered accurate if its $R$geq 0.9, relatively accurate if 0.8 $\leq$ $R$ $<$ 0.9, and contaminated if 0.5 $\leq$ $R$ $<$ 0.8.
VI.1 Detection rate and SNR
@@
We employ the reported sources to represent the outcomes following the $-\mathcal{F}$-statistic-analysis, and confirmed sources to signify those reported sources with an $R$ value exceeding the threshold $R_{t}$. The detection rate (DR) is succinctly formulated by their numbers as follows:
$\text{DR}=\frac{N_{c}}{N_{r}},$ (24)
@@
where $N_{r}$ is the number of reported sources, and $N_{c}$ is the number of confirmed sources. The evolutionary trajectory of our methodology through the various stages of $-\mathcal{F}$-statistic-analysis (elaborated in Sec. V) is vividly depicted in the histogram of $R$ presented in Fig. 14. In the initial step, a significant reduction in individual degeneracy noise was achieved, resulting in 25,615 reported sources with a DR of 23.2% at $R_{t}=0.8$ (A). The second step concentrated on diminishing instrument noise through cross-validation with a secondary search, yielding 10,623 reported sources and a DR of 45.2% at $R_{t}=0.8$ (B2). Subsequent refinement in the third step, which confined Galactic Latitude within ±$\pm 0.5$ rad, led to a reduction to 9,143 reported sources and a DR of 52.0% at $R_{t}=0.8$ (B3). Finally, the fourth step addressed the difference in $\mathcal{F}$-statistic value and overlapping degeneracy noise with $\mathcal{F}$-statistic values surpassing those of any individual signal, resulting in 6,508 reported sources with 4112 confirmed sources ($R_{t}=0.8$) and DR = 63.2%.
@@
Figure 14: The histogram depicts the distribution of correlation coefficients in the results after each step. A, B2, B3, and C correspond to the four steps outlined in Sec. V. A signifies the outcomes after eliminating individual signal degeneracy noise (Sec. V.1). B2 represents the results after cross-validation with the second search (Sec. V.2.2). B3 indicates the outcomes after restricting the Galactic Latitude (Sec. V.2.3), and C denotes the reported sources after removing overlapping degeneracy noise (Sec. V.3).
@@
Owing to signal overlap, the $\mathcal{F}$-statistic value for each source may receive contributions from other sources, resulting in an inflated SNR. To illustrate this effect, we depict the SNR-ratio between confirmed sources ($R_{t}=0.8$) and their corresponding injection sources in Fig. 15.
[…]
@@
The sources are arranged in ascending frequency order, and to achieve a smoother curve, we take an average for every 50 sources. In Fig. 15, “unremove degeneracy noise” refers to the SNR before the $-\mathcal{F}$-statistic analysis, while “remove degeneracy noise” corresponds to the SNR after the $-\mathcal{F}$-statistic analysis (the steps for removing degeneracy noise are outlined in Sec. V.1 and Sec. V.3). The average SNR-ratio is 1.21 and 1.17 for “unremove degeneracy noise” and “remove degeneracy noise”, respectively. After the $-\mathcal{F}$-statistic analysis, the fact that the average SNR-ratio approaches 1 indicates a closer match between the SNR of the confirmed sources and their corresponding injection sources, as the $\mathcal{F}$-statistic component from other sources has been eliminated. As the frequency increases, the SNR-ratio gradually converges to 1. However, at lower frequencies ($f<2\times 10^{-3}$ Hz), a substantial number of unresolved Galactic binaries contribute energy (degeneracy noise) to the confirmed sources, leading to $\text{SNR}_{\text{{con}}}$ being significantly larger than their corresponding $\text{SNR}_{\text{{inj}}}$.